HUBBLE SPACE TELESCOPE DAILY REPORT #2219 PERIOD COVERED: 0000Z (UTC) 09/29/98 - 0000Z (UTC) 09/30/98 Daily Status Report as of 273/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD/MA2/WF/PC-2/NIC/1/2/3 8058 (HDF-South Observations) The Space Telescope Imaging Spectrograph (CCD and MA2), the WF/PC-2, and the Near Infrared Camera (Cameras One, Two and Three) were all used to make long exposure Hubble Deep Field (HDFS-22333) observations. This program is designed to provide deep observations of a high-galactic latitude field in the southern continuous viewing zone. The STIS field is centered on a bright QSO at z=2.25. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 12 Successful : 12 2.2 FHST Updates: Scheduled : 4 Successful : 4 2.3 Operations Notes: The Hubble Deep Field - South (HDF-S) campaign continues with no significant problems reported. The NSSC-1 status buffer was dumped and reset two times as directed by ROP NS-3, and as expected in the HDF-S timeframe. ROP NS-5 was invoked once to reset an SI C&DH error. NS-13 was used to perform an NSSC-1 telemetry diagnostic. As a result, at 272/1601Z the NSSC-1 limit for memory temperature 1 was raised. Using ROP SR-1, the SSR EDAC error counter was reset three times. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH